OpenStudy (anonymous):

find the initial value of \[y^2 dy/dt=y^2+ty-t^2\] . please help me

OpenStudy (anonymous):

oh i forgot y(e)=0

sam (.sam.):

Have you tried using y=vt?

OpenStudy (anonymous):

i havent tried yet, im trying now

OpenStudy (anonymous):

can you please show me how ? i've tried let y=vt and dy/dv=v+t dv/dt but still cant get it

sam (.sam.):

From y=vt, you take dy/dt not dy/dv

OpenStudy (anonymous):

opps sorry typo..

sam (.sam.):

Sub it in \[t^2v^2(v+t\frac{dv}{dt})=t^2v^2+t^2v-t^2\] Simplify/Factor \[t^2v^2(v+t\frac{dv}{dt})=t^2(v^2+v-1)\] \[\frac{dv }{dt}=\frac{-v^3+v^2+v-1}{v^2t}\]

sam (.sam.):

\[\int\limits \frac{v^2}{-v^3+v^2+v-1}dv=\int\limits \frac{1}{t}dt\]

sam (.sam.):

There you go

OpenStudy (anonymous):

omg, thank you so much

sam (.sam.):

yw